OpenStudy (anonymous):

What is 3 to the power of 3 over 2 equal to?

OpenStudy (jdoe0001):

well... what's \(\bf 3^3 \quad ?\)

OpenStudy (anonymous):

it is 27

OpenStudy (anonymous):

3x3x3 = 27

OpenStudy (jdoe0001):

well, then \(\bf \cfrac{3^3}{2}\implies \cfrac{27}{2}\implies 13\frac{1}{2}\implies 13.5\)

OpenStudy (anonymous):

no i mean like 3^3/2

OpenStudy (jdoe0001):

\(\bf {\color{blue}{ \cfrac{3^3}{2}}}\implies \cfrac{27}{2}\implies 13\frac{1}{2}\implies 13.5\)

OpenStudy (anonymous):

thats the same thing

OpenStudy (anonymous):

3^3/2

OpenStudy (anonymous):

that is my question

OpenStudy (jdoe0001):

ohhh I see Ithink you mean \(\huge 3^{\frac{3}{2}}\quad ?\)

OpenStudy (anonymous):

yes perfect

OpenStudy (jdoe0001):

\(\large \bf{ 3^{\frac{3}{2}} \\ \quad \\ a^{\frac{{\color{blue} n}}{{\color{red} m}}} = \sqrt[{\color{red} m}]{a^{\color{blue} n}} \qquad thus \\ \quad \\ 3^{\frac{3}{2}}\implies \sqrt[{\color{red}{ 2}}]{3^{\color{blue}{ 3}}} }\implies \sqrt{3\cdot 3\cdot 3}\implies \sqrt{3\cdot 3}\cdot \sqrt{3}\implies \sqrt{3^2}\cdot \sqrt{3}\) what would that simplify to?

OpenStudy (anonymous):

id really know...

OpenStudy (anonymous):

i am bad at this concept

OpenStudy (jdoe0001):

well, just keep in mind that, any FACTOR inside the radical, that matches the root, goes "over the root" fence per se =) so \(\large \bf \sqrt[{\color{blue}{ n}}]{x^{\color{blue}{ n}}} \iff x\) so.. which one do you think we could take out?

OpenStudy (anonymous):

i honestly dont know but if u want options, here: cube root of 9 square root of 9 cube root of 27 square root of 27

OpenStudy (anonymous):

i think it is the last one... what do u say?

OpenStudy (jdoe0001):

well \(\bf 3^{\frac{3}{2}}\implies \sqrt[{\color{red}{ 2}}]{3^{\color{blue}{ 3}}}\implies \sqrt{3\cdot 3\cdot 3}\implies \sqrt{27}\)

OpenStudy (anonymous):

so im right?

OpenStudy (jdoe0001):

yeap

OpenStudy (jdoe0001):

3*3*3 = 27